# Exploring language relations through syntactic distances and geographic proximity

### v1

@@ I Introduction
A particularly simple but elucidating approach to analyze syntactic variations is by means of parts of speech (POS). These denote word classes with well defined grammatical functions that share common morphological properties Manning and Schutze (1999). For instance, almost all languages distinguish between verbs and nouns, i.e., roughly between actions and entities. As a consequence, one can categorize words or lexical items as members of any of the proposed POS, typically around 15. This classification has its own limitations (e.g., certain languages do not distinguish between verbs and adjectives) but it has the advantage of simplicity while capturing at the same time a large amount of morphosyntactic information. The POS approach has been proven especially useful in natural language processing tasks. The reason is that POS reveal much not only about the syntactic category of a word but also about that of its neighboring words, due to semantic restrictions. For example, if a word is a noun it will most likely be surrounded by a determiners and adjectives, forming a noun phrase. Therefore, we can gain insight about the phrasal structure of a language by examining POS sequences. This is precisely the main objective of this work: to model these sequences as stochastic processes, analyze their correlations and compute syntactic distances between languages using POS sequence distributions taken from a multilingual corpus.
Statistics of POS sequences, specifically the analysis of POS $r$-grams, defined as sequences of $r$ successive POS, has served for diverse research goals. One may hypothesize that genres are characterized by different syntactic structures, which would then allow for reliable genre classification. As demonstrated in Ref. Feldman et al. (2009), a careful study of POS trigram histograms provides a high-performance genre classifier. Strikingly, series of POS trigrams can be employed for building phylogenetic language trees solely from translations Rabinovich et al. (2017). The premise here is that syntactic features are retained in the translation process. Furthermore, assuming that POS tags can be predicted for historically close languages it is possible to train a language model to measure proximity among languages Samohi et al. (2022). However, these previous works set the POS block length in a somewhat heuristic manner. Below, we demonstrate using information-theoretic methods that trigrams suffice to account for the correlations present in POS sequences. In other words, it is not necessary to consider $r$-grams with $r\geq 4$ to gain more information, a result that considerably simplify analyses that involve POS series.
@@
Interestingly, we find a correlation between the obtained linguistic distance and the geographic distance spanned between locations assigned to each language. These are WALS locations, which correspond to the geographic coordinates associated to the centre of the region where the analyzed languages are spoken Dryer and Haspelmath (2013). Despite the fact that the centres are obviously only approximate and that our calculation of the linguistic distances has its own limitations, we clearly find that most of the syntactically close languages are also geographically close. This is expected since similar language usually lie in a continuum but there exist conspicuous exceptions, as we shall discuss below.
@@ II.1 Data
Importantly, we select languages within a contiguous region (except for Afrikaans which is included since this belongs to the Indo-European family) because we shall later explore possible correlations between linguistic and geographic distances. Specifically, we focus on Eurasia since this is a single continent with both a rich linguistic diversity and abundant data availability.
@@ II.3 Estimation from finite data
$\begin{split}\hat{p}_{{}_{\mathcal{L}}}(z_{i_{r}}|z_{i_{1}},\ldots,z_{i_{r-1}})\equiv\hat{P}(X_{r}=z_{i_{r}}|X_{1}=z_{i_{1}},\ldots,X_{r-1}=z_{i_{r-1}})=\dfrac{\hat{n}^{(r)}_{(i_{1},\ldots,i_{u},i_{r})_{L}}}{\sum\limits_{v=0}^{L-1}\hat{n}^{(r)}_{(i_{1},\ldots,i_{r-1},v)_{L}}},\end{split}$ (10)
@@ II.4.1 Predictability gain
Our first objective is to quantify the information gained in POS sequences of varios languages when performing predictions considering $(u+1)$th-order transition probabilities rather than $u$th-order.
@@ II.5 Language distances
$\begin{split}d_{{}_{JS}}(\mathcal{L},\mathcal{L}^{\prime})&=\sqrt{\dfrac{1}{2}\sum_{j=0}^{L^{{}^{3}}-1}\left(p_{{}_{\mathcal{L}}}(b_{j})\log\left(\dfrac{2p_{{}_{\mathcal{L}}}(b_{j})}{p_{{}_{\mathcal{L}}}(b_{j})+p_{{}_{\mathcal{L}^{\prime}}}(b_{j})}\right)+p_{{}_{\mathcal{L}^{\prime}}}(b_{j})\log\left(\dfrac{2p_{{}_{\mathcal{L}^{\prime}}}(b_{j})}{p_{{}_{\mathcal{L}}}(b_{j})+p_{{}_{\mathcal{L}^{\prime}}}(b_{j})}\right)\right)}\,\end{split}$ (16)
@@
Following a similar methodology, for each language pair within our dataset we compute their corresponding JS distances using Eqs. (16), yielding a distance matrix of dimension $67\times 67$, which bases our subsequent clustering analysis. This methodology considers all available data for each language. An alternative approach is presented in App. D, where we calculate distances among text samples extracted from the same language group. We now present the results obtained with the JS distance. Importantly, we also obtain similar results employing a different metric (the Hellinger distance, see App. E), which reinforces the validity of our findings.
@@ III.1.1 Distance matrix
The distance matrix generated from the data with the aid of Eq. (16) can be better visualized through a clustermap (see Fig. 4). This representation employs hierarchical clustering Nielsen (2016) and heatmap visualization. We use the complete linkage method Defays (1977) for clustering, organizing rows and columns based on similarity, depicted as dendrograms in the same figure. The heatmap, which represents distance values by a color spectrum as indicated in the inset of Fig. 4, enables a comprehensive exploration of our data relationships and structure.
@@ III.1.2 Language tree and k-medoids clustering
First, we construct a network whose nodes represent languages whereas the edges’ weight are given by their pairwise distances. Only the edges that generate a minimum spanning tree are considered as part of our analysis. The minimum spanning tree is a structure that connects all the data points in our dataset results with the minimum possible total edge weight Gower and Ross (1969). We depict the tree in Fig. 5, employing the Kamada-Kawai layout Kamada et al. (1989). This is a force-directed layout, implying that, for weighted graphs as in our case, the length of the edges tend to be larger for heavier weights. This visualization provides an intuitive representation of the relationships encoded in the minimum spanning tree, offering insights into the overall structure and connectivity patterns, not only among closely related languages, but also between distinct language families. To enhance our interpretability, we assign consistent colors to items within the same cluster, obtained from the k-medoids analysis mentioned earlier. Nodes from the same language group are linked with a full line; nodes from the same family but different group are linked with a dashed line; finally, nodes belonging to distinct families are connected with dotted lines. The shape of the nodes represents the language type: circles are assigned to fusional languages, squares to the agglutinative type and diamonds to isolating languages. The languages plotted with half nodes are both isolating and another type. For example, English is plotted with a half diamond node because it is considered to be a fusional-isolating language Haselow (2011).
Figure 5: Minimum spanning tree generated from the Jensen-Shannon distance matrix, with node colors representing clusters identified through k-medoids analysis. Nodes’ shape represent the language typology. Full lines are assigned to links between languages belonging to the same group; dashed lines for languages of the same family but different group; finally, dotted lines connect languages from distinct families.
@@
Overall, our cluster results are consistent with well established families and linguistic groups. We indeed observe departures that could be attributed to methodology inaccuracies. However, another interpretation is possible. Very recently, a family tree based on the phylogenetic signal of syntactic data has been inferred Hartmann and Walkden (2024), pointing to salient deviations with respect to the trees derived from the comparative method, which typically did no take into account syntactic data. Therefore, language groups formed upon syntactic analyses need not to fully agree with those groups that emerge from phonetic or lexical similarities.
@@ IV Conclusions
In this work we have collected and analyzed POS tagged sentences available in the Universal Dependencies library for $67$ contemporary languages located in a geographically contiguous region. Following an information-theoretic approach, we have provided evidence showing the effectiveness of utilizing the trigram probability distribution of parts-of-speech for characterizing the syntax statistics of languages. Through this method, we have computed distances between languages by calculating statistical divergences between trigram distributions, revealing both well-established language groupings and less familiar but already documented linguistic connections. Furthermore, our analysis has delved into the correlation between linguistic and geographic distances. We have found that spatially proximate languages tend to exhibit more similar morphosyntactic characteristics compared to those located farther apart. The obtained results suggest a logarithmic relation between these distances. This finding is in fact in good agreement with results reported in Refs. Nerbonne (2010); Jäger (2018), where different measures of linguistic distances are defined.
@@
Looking forward, our research suggests several promising avenues for future exploration. Firstly, expanding the range of languages considered in our analysis, even including different varieties of certain languages, could provide a more comprehensive understanding of linguistic diversity and relationships across different language families. Additionally, investigating the evolution of morphosyntactic similarities over time holds great potential for uncovering insights into the dynamical nature of grammar and word formation systems and their historical interconnections. For that, however, one would need multilingual diachronic corpora with a sufficiently long time span.
V Data availability
The datasets of parts-of-speech for all languages analyzed during the current study are available in the Universal Dependencies repository. The dataset of the geographic location of each language is available in the WALS repository.
VI Acknowledgments
@@ Appendix A List of languages
We present in Table LABEL:tab:languages a list of the $67$ languages included in our study, categorized by language family, group, and morphological type.
@@ Appendix C Memory effects in POS sequences
Figure C.1: Estimated predictability gain when considering $(u+1)$th-order instead of $u$th-order transition probabilities in sequences of UPOS for (a) German, (b) Icelandic, (c) Portuguese and (d) Czech, shown in red. The mean and standard deviation of the predictability gain estimated on numerically generated sequences of memory $2$, generated from the estimated transition probabilities of each language, are represented in black.
@@ Appendix D Distance between texts belonging to a single language group
We consider a single linguistic group. For each language in this group we randomly select tagged sentences from our database until we reach approximately $10^{4}$ tokens. We iterate this procedure a maximum of $20$ times, without replacement (for a few languages there is not enough data to perform this procedure $20$ times). Then, for each text portion of each language inside the given group we calculate their probability distribution of POS trigrams and compute the pairwise distances among texts with the JS metric. In Fig. D.1 we present heatmaps corresponding to the distance matrices obtained for Germanic (panel a) and Slavic (panel b) texts. We can observe that, in general, the distance between texts of the same language is smaller when compared to texts coming from distinct languages. This holds even for languages that are considered to be closely related, such as Croatian and Serbian, and Belarusian, Russian and Ukrainian.
@@ Appendix E Clustermap obtained from the Hellinger distance matrix
Appendix F Silhoutte analysis and clusters obtained with k-medoids algorithm
@@
The clusters obtained using this method are presented in Table F.2.

### v2

@@ I Introduction
A particularly simple but elucidating approach to analyze syntactic variations is by means of parts of speech (POS). These denote word classes with well defined grammatical functions that share common morphological properties Manning and Schutze (1999). For instance, almost all languages distinguish between verbs and nouns, i.e., roughly between actions and entities. As a consequence, one can categorize words or lexical items as members of any of the proposed POS, typically around 15. This classification has its own limitations (e.g., certain languages do not distinguish between verbs and adjectives) but it has the advantage of simplicity while capturing at the same time a large amount of morphosyntactic information. The POS approach has been proven especially useful in natural language processing tasks. The reason is that POS reveal much not only about the syntactic category of a word but also about that of its neighboring words, due to semantic restrictions. For example, if a word is a noun it will most likely be surrounded by determiners and adjectives, forming a noun phrase. Therefore, we can gain insight about the phrasal structure of a language by examining POS sequences. This is precisely the main objective of this work: to model these sequences as stochastic processes, analyze their correlations and compute syntactic distances between languages using POS sequence distributions taken from a multilingual corpus.
Statistics of POS sequences, specifically the analysis of POS $r$-grams, defined as sequences of $r$ successive POS, has served for diverse research goals. One may hypothesize that genres are characterized by different syntactic structures, which would then allow for reliable genre classification. As demonstrated in Ref. Feldman et al. (2009), a careful study of POS trigram histograms provides a high-performance genre classifier. Strikingly, series of POS trigrams can be employed for building phylogenetic language trees solely from translations Rabinovich et al. (2017). The premise here is that syntactic features are retained in the translation process. Furthermore, assuming that POS tags can be predicted for historically close languages it is possible to train a language model to measure proximity among languages Samohi et al. (2022). However, these previous works set the POS block length in a somewhat heuristic manner. Below, we demonstrate using information-theoretic methods that trigrams suffice to account for the correlations present in POS sequences. In other words, it is not necessary to consider $r$-grams with $r\geq 4$ to gain more information, a result that considerably simplifies analyses that involve POS series.
@@
Interestingly, we find a correlation between the obtained linguistic distance and the geographic distance spanned between locations assigned to each language. These are WALS locations, which generally correspond to the geographic coordinates associated to the centre of the region where the analyzed languages are spoken. However, for some languages, the regions where they are spoken are discontinuous. In such cases, the locations are placed within the larger region in which the language is spoken Dryer and Haspelmath (2013). Despite the fact that the centres are obviously only approximate and that our calculation of the linguistic distances has its own limitations, we clearly find that most of the syntactically close languages are also geographically close. This is expected since similar language usually lie in a continuum but there exist conspicuous exceptions, as we shall discuss below.
@@ II.1 Data
Importantly, we select languages within a contiguous region (except for Afrikaans which is included since it belongs to the Indo-European family) because we shall later explore possible correlations between linguistic and geographic distances. Specifically, we focus on Eurasia since this is a single continent with both a rich linguistic diversity and abundant data availability.
@@ II.3 Estimation from finite data
$\begin{split}\hat{p}_{{}_{\mathcal{L}}}(z_{i_{r}}|z_{i_{1}},\ldots,z_{i_{r-1}})\equiv\hat{P}(X_{r}=z_{i_{r}}|X_{1}=z_{i_{1}},\ldots,X_{r-1}=z_{i_{r-1}})=\dfrac{\hat{n}^{(r)}_{(i_{1},\ldots,i_{r-1},i_{r})_{L}}}{\sum\limits_{v=0}^{L-1}\hat{n}^{(r)}_{(i_{1},\ldots,i_{r-1},v)_{L}}},\end{split}$ (10)
@@ II.4.1 Predictability gain
Our first objective is to quantify the information gained in POS sequences of various languages when performing predictions considering $(u+1)$th-order transition probabilities rather than $u$th-order.
@@ II.5 Language distances
$\begin{split}d_{{}_{JS}}(\mathcal{L},\mathcal{L}^{\prime})&=\\ &\sqrt{\dfrac{1}{2}\sum_{j=0}^{L^{{}^{3}}-1}\left(p_{{}_{\mathcal{L}}}(b_{j})\log\left(\dfrac{2p_{{}_{\mathcal{L}}}(b_{j})}{p_{{}_{\mathcal{L}}}(b_{j})+p_{{}_{\mathcal{L}^{\prime}}}(b_{j})}\right)+p_{{}_{\mathcal{L}^{\prime}}}(b_{j})\log\left(\dfrac{2p_{{}_{\mathcal{L}^{\prime}}}(b_{j})}{p_{{}_{\mathcal{L}}}(b_{j})+p_{{}_{\mathcal{L}^{\prime}}}(b_{j})}\right)\right)}\,.\end{split}$ (16)
@@
Following a similar methodology, for each language pair within our dataset we compute their corresponding JS distances using Eqs. (16), yielding a distance matrix of dimension $67\times 67$, on which our subsequent clustering analysis is based. This methodology considers all available data for each language. An alternative approach is presented in App. D, where we calculate distances among text samples extracted from the same language group. We now present the results obtained with the JS distance. Importantly, we also obtain similar results employing a different metric (the Hellinger distance, see App. E), which reinforces the validity of our findings.
@@ III.1.1 Distance matrix
The distance matrix generated from the data with the aid of Eq. (16) can be better visualized through a clustermap (see Fig. 4). This representation employs hierarchical clustering Nielsen (2016) and heatmap visualization. We use the complete linkage method Defays (1977) for clustering, organizing rows and columns based on similarity, depicted as dendrograms in the same figure. The heatmap, which represents distance values by a color spectrum as indicated in Fig. 4, enables a comprehensive exploration of our data relationships and structure.
@@ III.1.2 Language tree and $k$-medoids clustering
First, we construct a network whose nodes represent languages whereas the edges’ weight are given by their pairwise distances. Only the edges that generate a minimum spanning tree are considered as part of our analysis. The minimum spanning tree is a structure that connects all the data points in our dataset results with the minimum possible total edge weight Gower and Ross (1969). We depict the tree in Fig. 5, employing the Kamada-Kawai layout Kamada et al. (1989). This is a force-directed layout, implying that, for weighted graphs as in our case, the length of the edges tends to be larger for heavier weights. This visualization provides an intuitive representation of the relationships encoded in the minimum spanning tree, offering insights into the overall structure and connectivity patterns, not only among closely related languages, but also between distinct language families. To enhance our interpretability, we assign consistent colors to items within the same cluster, obtained from the $k$-medoids analysis mentioned earlier. Nodes from the same language group are linked with a full line; nodes from the same family but different group are linked with a dashed line; finally, nodes belonging to distinct families are connected with dotted lines. The shape of the nodes represents the language type: circles are assigned to fusional languages, squares to the agglutinative type and diamonds to isolating languages. The languages plotted with half nodes are both isolating and another type. For example, English is plotted with a half diamond node because it is considered to be a fusional-isolating language Haselow (2011).
Figure 5: Minimum spanning tree generated from the Jensen-Shannon distance matrix, with node colors representing clusters identified through $k$-medoids analysis. The shape of the nodes represents the language typology. Full lines are assigned to links between languages belonging to the same group; dashed lines for languages of the same family but different group; finally, dotted lines connect languages from distinct families.
@@
Overall, our cluster results are consistent with well established families and linguistic groups. We indeed observe departures that could be attributed to methodology inaccuracies. However, another interpretation is possible. Very recently, a family tree based on the phylogenetic signal of syntactic data has been inferred Hartmann and Walkden (2024), pointing to salient deviations with respect to the trees derived from the comparative method, which typically did not take into account syntactic data. Therefore, language groups formed upon syntactic analyses need not fully agree with those groups that emerge from phonetic or lexical similarities.
In App. G, we present the results of a similar analysis to the one conducted in this section using tetragrams instead of trigrams. The findings from both methods are consistent, providing strong evidence for our claim that the probability distribution of POS trigrams is sufficient to capture syntactic information in the analyzed languages.
@@ IV Conclusions
In this work we have collected and analyzed parts of speech tagged sentences available in the Universal Dependencies library for $67$ contemporary languages located in a geographically contiguous region. Following an information-theoretic approach, we have provided evidence showing the effectiveness of utilizing the trigram probability distribution of parts-of-speech for characterizing the syntax statistics of languages. Through this method, we have computed distances between languages by calculating statistical divergences between trigram distributions, revealing both well established language groupings and less familiar but already documented linguistic connections. Whereas most analyses of language families are conducted at the phonetic level, our syntactic approach, while yielding similar results, points to a robustness of linguistic family classifications across different linguistic levels. This opens up new avenues for linguistic research where syntactic data can complement phonetic data, providing a more comprehensive understanding of language evolution, classification and geographic relationships. We also stress that our approach is synchronic and thus complements the diachronic views more often encountered in historical linguistics.
A potential impact of our work is in the field of language documentation and revitalization. By quantifying syntactic similarities, our method can help identify which languages are most similar to endangered languages, therefore guiding efforts to develop educational materials and resources. For instance, if an endangered language has not been extensively documented, educational content from syntactically similar languages can be adapted more efficiently, preserving linguistic heritage more effectively. Furthermore, our approach can be instrumental in computational linguistics, particularly in the development of multilingual natural language processing systems. Understanding syntactic similarities between languages can improve the performance of machine translation systems, especially for low-resource languages. Our results suggest that POS trigrams capture essential syntactic structures, which can be easily integrated into algorithms to enhance cross-linguistic transfer learning. As another potential application, we mention that knowing the language distances across different levels can be useful in language teaching, as it may help focus on the aspects that differ most between the student’s native language and the target language.
Furthermore, our analysis has delved into the correlation between linguistic and geographic distances. We have found that spatially proximate languages tend to exhibit more similar morphosyntactic characteristics compared to those located farther apart. Our results suggest a logarithmic relation between these distances. This finding is in fact in good agreement with results reported in Refs. Nerbonne (2010); Jäger (2018), where different measures of linguistic distances are defined. Due to the limitations of selecting a single point to represent the coordinates of an entire language, the results obtained can potentially be improved by considering more accurate methods to assign these locations, such as systematically selecting the regions where the languages are predominantly spoken, or by considering regional varieties that take into account the spatial variation of languages.
@@
A limitation of our research is that the UD dataset predominantly covers well documented languages. Therefore, less studied languages and dialects are underrepresented, which may limit the generalizability of our findings to the entire spectrum of global languages, since our results might not fully capture the syntactic diversity present in linguistic minorities. However, the UD library is frequently updated, regularly expanding existing corpora and incorporating new languages. It would be straightforward to extend our analysis to include more languages in future works.
Additionally, the use of POS tags, while effective for our analysis, might oversimplify the complexities of syntactic structures across different languages. POS tagging neglects many linguistic nuances and syntactic intricacies, which may lead to a loss of detailed information. This simplification can potentially affect the accuracy of our distance measurements, as some syntactic phenomena unique to specific languages may not be adequately represented. For example, languages with rich morphological systems or unique syntactic constructions may have their complexities reduced to basic categories that do not fully reflect their syntactic richness. Future work should consider integrating more detailed morphosyntactic annotations, such as dependency and constituency parses, which are already included in the UD library, to capture a broader range of syntactic features.
@@ Appendix A List of languages
We present in Table LABEL:tab:languages a list of the $67$ languages included in our study, categorized by language family, group, and morphological type. The latter is a simplified approximation that is included here for completeness. An exhaustive morphological analysis of the explored languages is beyond the scope of the present work.
@@ Appendix C Memory effects in POS sequences
Figure C.1: Estimated predictability gain when considering $(u+1)$th-order instead of $u$th-order transition probabilities in sequences of POS for (a) German, (b) Icelandic, (c) Portuguese and (d) Czech, shown in red. The mean and standard deviation of the predictability gain estimated on numerically generated sequences of memory $2$, generated from the estimated transition probabilities of each language, are represented in black.
@@ Appendix D Distance between texts belonging to a single language group
We consider a single linguistic group. For each language in this group we randomly select tagged sentences from our database until we reach approximately $10^{4}$ tokens. We iterate this procedure a maximum of $20$ times, without replacement (for a few languages there is not enough data to perform this procedure $20$ times). Then, for each text portion of each language inside the given group we calculate its probability distribution of POS trigrams and compute the pairwise distances among texts with the JS metric. In Fig. D.1 we present heatmaps corresponding to the distance matrices obtained for Germanic (panel a) and Slavic (panel b) texts. We can observe that, in general, the distance between texts of the same language is smaller when compared to texts coming from distinct languages. This holds even for languages that are considered to be closely related, such as Croatian and Serbian, and Belarusian, Russian and Ukrainian.
@@ Appendix E Clustermap obtained from the Hellinger distance matrix
Appendix F Silhouette analysis and clusters obtained with $k$-medoids algorithm
@@
The clusters obtained using this method are presented in Table LABEL:tab:clusters.
@@ 30 Greek
Appendix G Minimum spanning tree with tetragrams
Figure G.1: Minimum spanning tree generated from the Jensen-Shannon distance matrix, obtained from the probability distribution of POS tetragrams, with node colors representing clusters identified through $k$-medoids analysis. The shape of the nodes represents the language typology. Full lines are assigned to links between languages belonging to the same group; dashed lines for languages of the same family but different group; finally, dotted lines connect languages from distinct families.
In order to validate the results obtained in Sec. III.1.2 using trigrams, we perform a similar analysis using higher order $r$-grams for the $67$ selected languages.
Specifically, in Fig. G.1 we depict the minimum spanning tree obtained using tetragrams, which can be compared with the tree shown in Fig. 5. For tetragrams, we find that the optimal number of clusters is $32$ (instead of $31$ for trigrams). We observe that the obtained clusters and connections in the tree are very similar in both cases.
One noticeable difference is that in Fig. G.1 there is a connection between Arabic and Hebrew that was missing in the tree of Fig. 5. This is a nice finding since both languages are Semitic, although the $3$ languages of this family appear in $3$ different clusters. This and other discrepancies can be attributed to small data samples, which can contain low values of tokens (of the order of ten thousand for certain languages), whereas the number of possible tetragrams is $5062515^{4}=$ 50625.
Choosing the size of the $r$-grams requires a compromise between the complexity of syntactic structures considered and the accuracy of the parameters estimation. In Sec. II.4 we have provided evidence that the choice of trigrams offers a good solution to this compromise and the analysis performed with tetragrams supports this claim, since there exist no significant differences with the results based on trigrams.
Availability of data and material
The datasets of parts-of-speech for all languages analyzed during the current study are available in the Universal Dependencies repository at http://hdl.handle.net/11234/1-5287. The dataset of the geographic location of each language is available in the WALS repository at https://wals.info/languoid.
Acknowledgements
We thank S. J. Greenhill for useful discussions.
Funding